Sell or Process Further

Rise N’ Shine Coffee Company produces Columbian coffee in batches of 5,200 pounds. The standard quantity of materials required in the process is 5,200 pounds, which cost $5 per pound. Columbian coffee can be sold without further processing for $8.4 per pound. Columbian coffee can also be processed further to yield Decaf Columbian, which can be sold for $10 per pound. The processing into Decaf Columbian requires additional processing costs of $9,650 per batch. The additional processing will also cause a 6% loss of product due to evaporation.

a. Prepare a differential analysis dated October 6 on whether to sell regular Columbian (Alternative 1) or process further into Decaf Columbian (Alternative 2).

Differential Analysis
Sell Regular Columbian (Alt. 1) or Process Further into Decaf Columbian (Alt. 2)
October 6
Sell Regular
Columbian
(Alternative 1)
Process Further
into Decaf Columbian
(Alternative 2)
Differential
Effect on Income
(Alternative 2)
Revenues $ $ $
Costs
Income (Loss) $ $ $

b. Should Rise N’ Shine sell Columbian coffee or process further and sell Decaf Columbian?

c. Determine the price of Decaf Columbian that would cause neither an advantage nor a disadvantage for processing further and selling Decaf Columbian. Round your answer to two decimal places.
$ per pound
Answer #1

Solution a:

Differential Analysis
Sell Regular Columbian (Alt. 1) or Process Further into Decaf Columbian (Alt. 2)
Oct-06
Sell Regular Columbian Process Further into Decaf Columbian Differential Effect on Income
(Alternative 1) (Alternative 2) (Alternative 2)
Revenues 43680 48880 5200
Costs 9650 -9650
Income (Loss) 43680 39230 -4450

2. General - - Home Insert Page Layout Excel Poo-1.1 - Microsoft Excel - 7 X File Formulas Dala Review View $ 1 Calculate Now

Solution b:

Rise N’ Shine should sell Columbian coffee.

Solution c:

Loss to be recovered = $4450

Units of Decaf Columbian = 5200 *94% = 4888 units

New Price should be = $10 + ($4450/ 4888) = $10.91 per pound